OpenStudy (anonymous):

A fifth-degree polynomial factored over the integers is: p(x)=3x(x-2)(5x-7)(x^2-4x+8) How many zeros are real? How many zeros are non real?

OpenStudy (kinggeorge):

You know that at least 3 zeroes are real since you've factored the polynomial into 3 linear terms and 1 quadratic. The linear terms each represent a single real zero. This means we need to find how many real/complex roots the quadratic has.

OpenStudy (kinggeorge):

To do this, you'll want to look at the discriminant \(b^2-4ac\). If we calculate this, we get\[b^2-4ac=16-4(8)=-16\]Since this is negative, we must have two complex zeroes.

OpenStudy (anonymous):

what is the difference between a real zero and a complex zero?

OpenStudy (kinggeorge):

A complex zero is a zero that looks like the form \(a+bi\) where \(i=\sqrt{-1}\)

OpenStudy (anonymous):

so it is not real?

OpenStudy (kinggeorge):

correct

OpenStudy (anonymous):

so that means it has 3 real zeros and 2 non real zeros?

OpenStudy (kinggeorge):

That is correct.

OpenStudy (anonymous):

thanks

OpenStudy (kinggeorge):

You're welcome.